What is the probability that a pet owning household does not own a fish when the total is 69 and there is 6.2 that do own fish
Answer by stanbon(57219)

What is the probability that a pet owning household does not own a fish when the total is 69 and there is 6.2 that do own fish
---
P(do own fish) = 6.2/69
---
P(do NOT own fish) = 1 - (6.2/69) = 0.9101
=========
Cheers,
Stan H.
=========